 Use the inductive reasoning to predict the next term in each list.


1. 4, 8, 12, 16, 20, 24, ?


2. 5, 11, 17, 23, 29, 35, ?


3. 1, 8, 27, 64, 125, ?


4. -2, 3, -4, 5, -6, 7, ?


5. 2, 5, 10, 17, 26, ?





1
Expert's answer
2022-01-06T16:45:53-0500

1) "a_{n+1}=a_n+4."

"a_7=a_6+4=24+4=28."


2) "a_{n+1}=a_n+6."

"a_7=a_6+6=35+6=41."


3) "a_n=n^3."

"a_6=6^3=216."


4) "a_{n+2}=a_n-2."

"a_7=a_5-2=-6-2=-8."


5) "a_{n+1}=a_n+2n+1."

"a_6=a_5+2*5+1=26+11=37."